hello tutors, by the i'd like to ask a question about problem solving. 'cause i have no idea of this problem can you help me to solve this problem. the problem is this...
a mother is 36 years old and her daughter is 14 years old. in how many years will the mother be twice as old as the daughter?
thanks you guy's!.. hope you can help me with my homework.

Answer by ankor@dixie-net.com(22740)

a mother is 36 years old and her daughter is 14 years old. in how many years will the mother be twice as old as the daughter?
:
let y = no. of years for this to be accomplished
:
36 + y = 2(14 + y)
:
36 + y = 28 + 2y
:
36 - 28 = 2y - y
:
y = 8 yrs for Mom to be twice as old as daughter
:
:
Check this:
36 + 8 = 2(14 + 8)
:
44 = 2(22)